### v1

@@ 3 Privacy-preserving community detection for locally distributed networks
This section proposes the privacy-preserving community detection algorithm for locally distributed networks. Before that, we specify the centralized estimator and develop a bias-adjustment procedure to remove the bias inherent in the RR-perturbed networks.
Notes and notation. We use $\lambda_{\min}(A)$ to denote the smallest non-zero eigenvalue of a given matrix $A$. Moreover, we use the projection distance $\mathrm{dist}(U,U^{\prime}):=\|UU^{\intercal}-U^{\prime}U^{\prime{\intercal}}\|_{2}$ to measure the subspace distance between any pair of orthogonal basis $U,U^{\prime}\in\mathbb{R}^{n\times K}$. $\|A\|_{2}$ denotes the spectral norm of matrix $A$ or the Euclidian norm of vector $A$ and $\|A\|_{\tiny F}$ denote the Frobenius norm of matrix $A$. In addition, we use the following standard notation for asymptotics. We write $f(n)\asymp g(n)$ or $f(n)=\Theta(g(n))$ if $cg(n)\leq f(n)\leq Cg(n)$ for some constants $0<c<C<\infty$, $f(n)\lesssim g(n)$ or $f(n)=O(g(n))$ if $f(n)\leq Cg(n)$ for some constant $C<\infty$, and $f(n)\gtrsim g(n)$ or $f(n)=\Omega(g(n))$ if $f(n)\geq cg(n)$ for some constant $c>0$. Finally, the constants throughout this work may be different from place to place.
Recall that there are $L$ local machines, with each owning a symmetric binary network adjacency matrix $A_{l},l=1,...,L$. $A_{l}$’s are assumed to be generated from a multi-layer SBM. Denote the population of $A_{l}$ by $P_{l}:=\Theta B_{l}\Theta^{\intercal}$, where $\Theta\in\mathbb{R}^{n\times K}$ is the membership matrix which is formed by forcing $\Theta_{ig_{i}}=1$ and $\Theta_{ij}=0$ for $j\neq g_{i}$. We aim to recover the common membership matrix $\Theta$ using the network information.
Without the privacy constraint, a common centralized approach is to perform community detection algorithms on $\sum_{l}A_{l}$ (Bhattacharyya and Chatterjee, 2018; Paul and Chen, 2020). However, direct summation of $A_{l}$’s would lead to information loss when $A_{l}$’s are heterogeneous in the sense of community pattern. In other words, if some $P_{l}$’s are positive semidefinite, which present homophily social ties and others are negative semidefinite, which indicate heterophily social ties (Kim and Leskovec, 2012), then adding them up would cancel out some community signals. To alleviate this, we consider the matrix
$M:=\frac{1}{Ln}\sum_{l=1}^{L}A_{l}^{2}$ (5)
studied in Lei and Lin (2022), which could accommodate different community patterns. Note that Lei and Lin (2022) further developed a bias-correction procedure to modify the diagonal elements of $M$. The centralized estimator is then obtained by performing spectral clustering on $M$. Specifically, the eigen-decomposition is first computed on $M$ to obtain the first $K$ eigenvectors, and then $k$-means is conducted on the resulting eigenvectors to obtain $K$ communities. Hopefully, $M$ can approximate its population counterpart
${Q}:=\frac{1}{L}\sum_{l=1}^{L}\frac{P_{l}^{2}}{n}:=\frac{1}{L}\sum_{l=1}^{L}{Q}_{l},$ (6)
whose eigenvector reveals the true communities in the sense that two rows of the eigenvector matrix are identical if and only if the corresponding nodes are in the same underlying communities (Rohe et al., 2011; Lei and Rinaldo, 2015; Zhang et al., 2022).
However, with the privacy constraint, we could only use the RR-perturbed adjacency matrix instead of $A_{l}$ in (5), which is a biased estimator of $P_{l}$. In the next subsection, we will develop a bias-adjustment procedure to obtain a good estimator for $Q$.
@@ 3.1 Bias-adjustment
Our debiasing procedure involves two steps: in step 1, we build a debiased estimator for the link probability matrix upon the RR-perturbed adjacency matrix stored in each local machine; in step 2, we further remove the bias of the diagonals caused by taking the square of the debiased estimated matrix obtained in step 1.
Step 1. We now illustrate the necessity of bias-adjustment when dealing with RR-perturbed data.
@@
and ${\tilde{A}}_{l,ji}={\tilde{A}}_{l,ij}$. By (7) and (8) and taking expectations for two times, we have
@@
which is biased from the true expectation $P_{l,ij}$ of $A_{l,ij}$. Therefore, for $i\neq j$, the local machine only needs to debias ${\tilde{A}}_{l,ij}$ to be
@@
Step 2. Though $\overline{A}_{l}$ is the unbiased estimator for $P_{l}$, the diagonal elements of $\overline{A}_{l}^{2}$ can cause bias for $P_{l}^{2}$. It turns out that the error would be smaller if we further subtract $\frac{\sum_{j}{\tilde{A}}_{l,ij}(q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}$ from the $i$th diagonal elements of $\overline{A}_{l}^{2}$ provided that $\max_{i,j,l}P_{l,ij}=o(1)$. In particular, this debiasing step is crucial when $L$ is increasing. For the sake of saving space, the details are relegated to Appendix A.
Combining the two steps of debiasing, we eventually obtain the following estimator $\overline{M}$ for $Q$ defined in (6),
$\overline{M}:=\frac{1}{L}\sum_{l=1}^{L}\overline{M}_{l}$ (10)
with
@@
Therefore, with the privacy constraint, the ideal centralized estimator is to perform spectral clustering on the debiased estimator $\overline{M}$. Noting that $\overline{M}_{l}$’s are separable across different local machines. We show in the next subsection how we can approximate the eigenvector of $\overline{M}$ with the local computation.
@@ 3.2 Privacy-preserving distributed spectral clustering
Since DP enjoys the post-processing property, in principle, the local machines can send the dense debiased estimators $\overline{M}_{l}$’s to the central server where the spectral clustering is further performed on the aggregated matrix $\overline{M}$. However, both the communication and computational cost can be very high for sending the whole dense matrices $\overline{M}_{l}$’s and performing the aggregation of these matrices on the central server, especially when the matrix dimensions are large. Instead, we consider a more efficient distributed algorithm described as follows.
We adopt the one-shot averaging framework in Charisopoulos et al. (2021) for distributed computing performed on our debiased RR-perturbed network matrices. That is, on the $l$th machine, we first compute the eigen-decomposition of $\overline{M}_{l}$, and obtain the matrix $\hat{V}_{l}\in\mathbb{R}^{n\times K}$ that consists of the first $K$ eigenvectors, for all $l=1,...,L$. On the central server, we average all the local solutions. Note that direct averaging is inferior due to the orthogonal ambiguity of the problem, i.e., for any orthogonal matrix $Z$, $\hat{V}_{l}Z$ also forms a valid orthonormal basis of the subspaces formed by $\hat{V}_{l}$. Hence, $\hat{V}_{l}$’s are not guaranteed to be close to each other, even if their corresponding subspaces are close. In light of this, the local $\hat{V}_{l}$’s must be aligned with each other before averaging.
To avoid the orthogonal ambiguity of the problem, we use the Procrustes fixing strategy (Schönemann, 1966; Cape, 2020; Charisopoulos et al., 2021; Guo et al., 2021). For two matrices $V_{1},V_{2}$ to be aligned, we can map $V_{2}$ to $V_{1}$ using an orthogonal transformation solved by the following optimization,
@@
where $\mathcal{O}_{k}$ denotes the set of $K\times K$ orthogonal matrices. Note that (12) has a closed-form solution. Let the SVD of $V_{2}^{\intercal}V_{1}$ be $LDR^{\intercal}$, then $Z:=LR^{\intercal}$ is the solution to (12). In such a way, we can align local eigenspaces $\hat{V}_{l}$ to an arbitrarily chosen reference eigenspace, for example, $\hat{V}_{o}$, after which the aligned eigenspaces can then be averaged and orthogonalized.
The proposed privacy-preserving Distributed Spectral Clustering (ppDSC) algorithm is summarized in Algorithm 1. It should be noted that, in a real scenario, the computation on the local machine may be fulfilled by a single party or cooperated by different parties. For example, a party with data ownership may desire to obtain good utility while preserving the data’s privacy. Thus all the local steps are conducted within the party. It can also happen that the RR perturbation is conducted by the data collection department of a company, while the data analysis department could do the remaining bias-correction and eigen-decomposition steps.
@@
1:  Input: Network adjacency matrices $\{A_{l}\}_{l=1}^{L}$ owned by local machines, number of clusters $K$, privacy parameters $q,q^{\prime}$, index of reference solution $o$.
2:  The local machines do:
3:  for $l=1$ to $L$ do
4:     RR-perturbation: Perturb $A_{l}$ to obtain ${\tilde{A}}_{l}$ using (7) and (8);
5:     Bias-correction: Constructs the debiased estimator $\overline{M}_{l}$ according to (11);
6:     Eigen-decomposition: Performs eigen-decomposition on $\overline{M}_{l}$ and denote the resulting eigen-spaces as $\hat{V}_{l}$.
7:  end for
8:  The central server does:
9:  Procrustes fixing: $\tilde{V}_{l}:=\hat{V}_{l}Z_{l}$ with $Z_{l}:=\operatorname{arg\,min}_{Z\in\mathcal{O}_{K}}\|\hat{V}_{l}Z-\hat{V}_{o}\|_{F}$ for each $l=1,...,L$;
10:  Aggregation: Aggregates the local solutions according to $\bar{V}:=\frac{1}{L}\sum_{l=1}^{L}\tilde{V}_{l}$;
11:  Orthogonalization: Performs the QR decomposition to obtain $\tilde{V},\tilde{R}={\rm qr}(\bar{V})$.
12:  Clustering: Conduct $k$-means clustering on $\tilde{V}$ with cluster number being $K$. The resulting clusters are formulated by the membership matrix $\tilde{\Theta}$.
13:  Output: The approximated eigen-space $\tilde{V}$ and the estimated membership matrix $\tilde{\Theta}$.
@@ 4 Theoretical analysis
To fix the ideas, we now recall and introduce some notations. We assume that there are $L$ local networks whose adjacency matrices $A_{l}\in\{0,1\}^{n\times n}$’s are generated from SBMs with layer-specific link probability matrices $B_{l}$’s and a consensus membership matrix $\Theta\in\mathbb{R}^{n\times K}$. We here assume that the number of communities $K$ is known. When $K$ is unknown, we will investigate the method and theory for the estimation of $K$, which is a model selection problem, as a future research topic, as this current work focuses on estimation and clustering analysis for a given model based on RR-perturbed network matrices. The number of nodes within community $k$ is denoted by $n_{k}$. The privacy parameters are denoted by $q,q^{\prime}$ ($1/2<q,q^{\prime}\leq 1$); see also (7) and (8). Our goal is to use ppDSC (Algorithm 1) to recover the first $K$ eigenvectors $V$ of the underlying target matrix $Q$ defined in (6) and the membership matrix $\Theta$.
The following assumptions are generally required to derive the theoretical results for community detection.
@@
The community is balanced such that for any $k\in\{1,...,K\}$,
$c\frac{n}{K}\leq n_{k}\leq C\frac{n}{K},$
for some constants $0\leq c\leq C$.
@@
The link probability matrix $B_{l}=\rho B_{l,0}$, where $\rho=o(1)$ and $B_{l,0}$ is a $K\times K$ symmetric matrix with constant entries in $[0,1]$.
@@
Assumption 4 (Limited heterogeneity)
The heterogeneity among the link probability matrices $B_{l}$’s are limited as
$\max_{l}\|B_{l}\Delta^{2}B_{l}-\frac{1}{L}\sum_{l=1}^{L}B_{l}\Delta^{2}B_{l}\|_{2}\leq\mathcal{H}(n,\rho,L).$
with $\Delta:={\operatorname{diag}}(\sqrt{n_{1}},...,\sqrt{n_{K}})$.
@@
Assumption 1 requires that the communities are balanced. Assumption 2 specifies the link probability matrices. In particular, we require that the network sparsity $\rho=o(1)$. Assumption 3 requires the linear growth of the aggregated squared link probability matrices in terms of its minimum eigenvalue, which is trivially satisfied when $B_{l,0}$’s are equal and is also used in Lei and Lin (2022). Assumption 3 combining with Assumptions 2 and 1 implies that $\sum_{l=1}^{L}B_{l}\Delta^{2}B_{l}$ is of full rank, and $Q$ is of rank $K$; see the proof of Corollary 1. Indeed, the full rank requirement could be extended to the rank-deficient case; see Zhang et al. (2022) for example. Assumption 4 bounds the maximum deviation of $B_{l}\Delta^{2}B_{l}$’s from its average by $\mathcal{H}(n,\rho,L)$. For simplicity, we call $\mathcal{H}$ the measure for the heterogeneity of locally stored networks.
@@
Suppose Assumptions 1 and 4 hold, the output $\tilde{V}$ of Algorithm 1 satisfies
$\mathrm{dist}(\tilde{V},V)\lesssim{\max_{l}\frac{\|\overline{M}_{l}-Q_{l}\|_{2}^{2}}{\lambda_{{\min}}^{2}(Q)}}+{\frac{\mathcal{H}^{2}(n,\rho,L)}{\lambda_{{\min}}^{2}(Q)}}+\frac{\|\sum_{l}\overline{M}_{l}/L-Q\|_{2}}{\lambda_{{\min}}(Q)}:=\mathcal{I}_{1}+\mathcal{I}_{2}+\mathcal{I}_{3}.$
@@
The condition that $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are smaller than a small constant $c$ is a standard condition (Damle and Sun, 2020; Charisopoulos et al., 2021); it ensures that the maximum local error $\max\|\overline{M}_{l}-Q\|_{2}$ is smaller than the eigen-gap $\lambda_{{\min}}(Q)$, which further indicates that the $K$-largest eigenvalues of $\overline{M}_{l}$ are separated from its remaining eigenvalues by Weyl’s inequality. In Corollary 1, we show that this condition is satisfied by our setting under mild conditions and further provide the explicit expressions for $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$.
@@
The decomposed three components $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$ are explained as follows. $\mathcal{I}_{1}$ corresponds to the maximum squared error of the local estimators $\hat{V}_{l}$’s except that we use $Q$ in the denominator instead of $Q_{l}$; see the Davis-Kahan theorem on the perturbation of eigenvectors (Davis and Kahan (1970); see Lemma 3 in the Appendix). As a result, we only require the averaged population matrix $Q:=\sum Q_{l}/L$ be of rank $K$, and the local population matrix $Q_{l}$ can be rank deficient, which is one of the benefits of combining the local eigenvectors. $\mathcal{I}_{2}$ corresponds to the degree of heterogeneity. To distinguish, we call $\mathcal{H}$ the measure for heterogeneity and call $\mathcal{I}_{2}$ the measure for scaled heterogeneity. It is worth noting that the scaled heterogeneity $\mathcal{I}_{2}$ affects the theoretical accuracy of $\tilde{V}$ according to its magnitude compared to $\mathcal{I}_{1}$ and $\mathcal{I}_{3}$. Moreover, $\mathcal{I}_{3}$ corresponds to the error of the centralized estimator $\hat{V}:=\frac{1}{L}\sum_{l}\overline{M}_{l}$. $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ can be regarded as the price needed to pay for the distributed computation.
One of the contributions in Theorem 2 is that it explicitly explains how the heterogeneity $\mathcal{H}$’s effect could be a curse or blessing to $\mathcal{I}_{2}$ and possibly the final error bound. Note that $\mathcal{H}$ and $\lambda_{{\min}}(Q)$ jointly decide $\mathcal{I}_{2}$. Therefore, if the local $B_{l}$’s are noisy from the signal such that combining them could hide the true signal, i.e., $\mathcal{H}$ is large, and $\lambda_{{\min}}(Q)$ is small, then the heterogeneity would be a curse. On the other hand, if the local $B_{l}$’s are far from the signal but combining them would enhance the signal, i.e., $\mathcal{H}$ is large, but $\lambda_{{\min}}(Q)$ can also be large such that their ratio is small, then the heterogeneity would be a blessing. We refer to the two simulated examples in Appendix C for more illustrations on the aforementioned phenomena.
To make the bounds in Theorem 2 more specific, in the following we provide Theorem 3 which establishes the estimation bounds of $\|\sum_{l}\overline{M}_{l}/L-Q\|_{2}$. Note that we assume that the privacy parameters $q,q^{\prime}$ relate to the privacy budget $\epsilon$ via (4). For illustration, we interchange the use of $q,q^{\prime}$, and $\epsilon$. Note that our results also hold for $q=\frac{\mathrm{e}^{\epsilon_{1}}}{1+\mathrm{e}^{\epsilon_{1}}}$ and $q^{\prime}=\frac{\mathrm{e}^{\epsilon_{2}}}{1+\mathrm{e}^{\epsilon_{2}}}$ with $\epsilon:=\max\{\epsilon_{1},\epsilon_{2}\}$ and $\epsilon_{1}$ and $\epsilon_{2}$ being of the same order.
@@
$\small\|\overline{M}-Q\|_{2}\leq C\max$$\bigg{\{}\frac{(q\rho+1-q^{\prime})^{1/2}n^{1/2}\rho\log^{1/2}(L+n)}{(q+q^{\prime}-1)\sqrt{L}},$
$\quad\quad\quad\frac{(q\rho+1-q^{\prime})\log(L+n)}{(q+q^{\prime}-1)^{2}\sqrt{L}},\,\frac{((q+q^{\prime}-1)\rho+1-q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}\bigg{\}},$ (14)

### v2

@@ 3 Privacy-preserving community detection for locally distributed networks
This section proposes the privacy-preserving community detection algorithm for locally distributed networks. In particular, we develop a bias-adjustment procedure to remove the bias inherent in the RR-perturbed networks.
Recall that there are $m$ local machines, and the $i$th machine owns $s$ adjacency matrices $A_{l\in S_{i}}$’s. $A_{l\in[L]}$’s are assumed to be generated from a multi-layer SBM. Denote the population of $A_{l}$ by $P_{l}:=\Theta B_{l}\Theta^{\intercal}$, where $\Theta\in\mathbb{R}^{n\times K}$ is the membership matrix which is formed by forcing $\Theta_{ig_{i}}=1$ and $\Theta_{ij}=0$ for $j\neq g_{i}$. We aim to recover the common membership matrix $\Theta$ using the network information.
Since $P_{l}$’s are not necessarily positive semidefinite, the top-$K$ singular vectors of $P_{l}$’s are a good choice to reveal the communities with both assortative and disassortative structures (Rohe et al., 2011), which is equivalent to the top-$K$ eigenvectors of $P_{l}^{2}$. In the following, we will develop a bias-adjustment procedure to obtain a good estimator for $P_{l}^{2}$ when we only have access to the RR-perturbed adjacency matrices.
@@ 3.1 Bias-adjustment
Our debiasing procedure involves two steps: in step 1, we build a debiased estimator for the connectivity matrix upon the RR-perturbed adjacency matrix stored in each local machine; in step 2, we further remove the bias of the diagonals caused by taking the square of the debiased estimated matrix obtained in step 1.
Step 1. We now illustrate the necessity of bias adjustment when dealing with RR-perturbed data.
@@
and ${\tilde{A}}_{l,ji}={\tilde{A}}_{l,ij}$. By (5) and (6), taking expectations for two times, we have
@@
which is biased from the true expectation $P_{l,ij}$ of $A_{l,ij}$. Therefore, for $i\neq j$, the local machine only needs to de-bias ${\tilde{A}}_{l,ij}$ to be
@@
Step 2. Though $\overline{A}_{l}$ is the unbiased estimator for $P_{l}$, the diagonal elements of $\overline{A}_{l}^{2}$ can cause bias for $P_{l}^{2}$. It turns out that the error would be smaller if we further subtract $\frac{\sum_{j}{\tilde{A}}_{l,ij}(q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}$ from the $i$th diagonal elements of $\overline{A}_{l}^{2}$ provided that $\max_{i,j,l}P_{l,ij}=o(1)$. In particular, this debiasing step is crucial when $L$ is increasing. To save space, the details are relegated to Section A of Supplementary Materials.
Combining the two steps of debiasing, we eventually obtain the following estimator $\overline{M}_{l}$ for $P_{l}^{2}/n$,
@@ 3.2 Privacy-preserving distributed spectral clustering
Since DP enjoys the post-processing property, in principle, each local machine can send the averaged estimator $\sum_{l\in S_{i}}\overline{M}_{l}/s$ to the central server where the spectral clustering is further performed on the aggregated matrix
$\overline{M}:=\frac{1}{m}\sum_{i=1}^{m}\sum_{l\in S_{i}}\frac{\overline{M}_{l}}{s}.$ (9)
However, both the communication and computational cost can be very high for sending the whole dense matrices $\sum_{l\in S_{i}}\overline{M}_{l}/s$’s and performing further aggregation of these matrices on the central server, especially when the matrix dimensions are large. Instead, we consider a more efficient distributed algorithm, as described below.
We adopt the one-shot averaging framework in Charisopoulos et al. (2021) for distributed computing performed on our debiased RR-perturbed network matrices. That is, on the $i$th machine, we first compute the eigen-decomposition on $\sum_{l\in S_{i}}\overline{M}_{l}/s$, and obtain the matrix $\hat{V}_{i}\in\mathbb{R}^{n\times K}$ that consists of the first $K$ eigenvectors, for all $i\in[m]$. On the central server, we average all the local solutions. Note that direct averaging is inferior due to the orthogonal ambiguity of the problem, i.e., for any orthogonal matrix $Z$, $\hat{V}_{i}Z$ also forms a valid orthonormal basis of the subspaces formed by $\hat{V}_{i}$. Hence, $\hat{V}_{i}$’s are not guaranteed to be close to each other, even if their corresponding subspaces are close. In light of this, the local $\hat{V}_{i}$’s must be aligned with each other before averaging.
To avoid the orthogonal ambiguity of the problem, we use the Procrustes fixing strategy (Schönemann, 1966; Cape, 2020; Charisopoulos et al., 2021; Guo et al., 2024). For two matrices $V_{1},V_{2}$ to be aligned, we can map $V_{2}$ to $V_{1}$ using an orthogonal transformation solved by the following optimization,
@@
where $\mathcal{O}_{k}$ denotes the set of $K\times K$ orthogonal matrices. Note that (10) has a closed-form solution. Let the SVD of $V_{2}^{\intercal}V_{1}$ be $LDR^{\intercal}$, then $Z:=LR^{\intercal}$ is the solution to (10). In such a way, we can align local eigenspaces $\hat{V}_{i}$ to an arbitrarily chosen reference eigenspace, for example, $\hat{V}_{o}$, after which the aligned eigenspaces can then be averaged and orthogonalized. As we will see in Section 4, the selection of the reference eigenspace does not affect the theoretical bounds.
1:  Input: Network adjacency matrices $\{A_{l}\}_{l=1}^{L}$, number of machines $m$, number of clusters $K$, privacy parameters $q,q^{\prime}$, index of reference solution $o$.
2:  The local machines do:
3:  for $i=1$ to $m$ do
4:     for $l$ in $S_{i}$ do
5:        RR-perturbation: Perturb $A_{l}$ to obtain ${\tilde{A}}_{l}$ using (5) and (6);
6:        Bias-correction: Constructs the debiased estimator $\overline{M}_{l}$ according to (8);
7:     end for
8:     Eigen-decomposition: Performs eigen-decomposition on $\sum_{l\in S_{i}}\overline{M}_{l}/s$ and denote the resulting eigen-spaces as $\hat{V}_{i}$.
9:  end for
10:  The central server does:
11:  Procrustes fixing: $\tilde{V}_{i}:=\hat{V}_{i}Z_{i}$ with $Z_{i}:=\operatorname{arg\,min}_{Z\in\mathcal{O}_{K}}\|\hat{V}_{i}Z-\hat{V}_{o}\|_{F}$ for each $i=1,...,m$;
12:  Aggregation: Aggregates the local solutions according to $\bar{V}:=\frac{1}{m}\sum_{i=1}^{m}\tilde{V}_{i}$;
13:  Orthogonalization: Performs the QR decomposition to obtain $\tilde{V},\tilde{R}={\rm qr}(\bar{V})$.
14:  Clustering: Conduct $k$-means clustering on $\tilde{V}$ with cluster number being $K$. The resulting clusters are formulated by the membership matrix $\tilde{\Theta}$.
15:  Output: The approximated eigen-space $\tilde{V}$ and the estimated membership matrix $\tilde{\Theta}$.
The proposed privacy-preserving Distributed Spectral Clustering (ppDSC) algorithm is summarized in Algorithm 1. It should be noted that, in a real scenario, the computation on the local machine may be fulfilled by a single party or cooperated by different parties. For example, a party with data ownership may desire to obtain good utility while preserving the data’s privacy. Thus, all the local steps are conducted within the party. It can also happen that the RR perturbation is conducted by the data collection department of a company, while the data analysis department could do the remaining bias-correction and eigen-decomposition steps.
@@ 4 Theoretical analysis
To fix the ideas, we now recall and introduce some notations. We assume that there are $L$ networks with adjacency matrices being $A_{l}\in\{0,1\}^{n\times n}$’s evenly distributed on $m$ local machines; each machine owns $s$ networks. For $i\in[m]$, the $i$th machine has networks indexed by set $S_{i}$. The adjacency matrices $A_{l}$’s are generated from SBMs with layer-specific connectivity matrices $B_{l}$’s and a consensus membership matrix $\Theta\in\mathbb{R}^{n\times K}$. We here assume that the number of communities $K$ is known. When $K$ is unknown, we will investigate the method and theory for the estimation of $K$, which is a model selection problem, as a future research topic, as this current work focuses on estimation and clustering analysis for a given model based on RR-perturbed network matrices. The number of nodes within community $k$ is denoted by $n_{k}$. The privacy parameters are denoted by $q,q^{\prime}$ ($1/2<q,q^{\prime}\leq 1$); see also (5) and (6). Our goal is to use ppDSC (Algorithm 1) to recover the first $K$ eigenspaces $V$ of the underlying target matrix $Q$ defined by
${Q}:=\frac{1}{L}\sum_{l=1}^{L}\frac{P_{l}^{2}}{n}:=\frac{1}{L}\sum_{l=1}^{L}{Q}_{l},$ (11)
and the membership matrix $\Theta$.
The following assumptions and definition are generally required to derive the theoretical results for community detection.
@@
The community is balanced such that for any $k\in\{1,...,K\}$, $c\frac{n}{K}\leq n_{k}\leq C\frac{n}{K}$ for some constants $0\leq c\leq C$.
@@
The connectivity matrix $B_{l}=\rho B_{l,0}$, where $\rho=o(1)$ and $B_{l,0}$ is a $K\times K$ symmetric matrix with constant entries in $[0,1]$.
@@
Definition 2 (Heterogeneity)
Suppose the connectivity matrix $B_{l}$’s are heterogeneous and define a heterogeneous measurement as
$\mathcal{H}(n,\rho,L):=\max_{i\in[m]}\|\frac{1}{s}\sum_{l\in S_{i}}B_{l}\Delta^{2}B_{l}-\frac{1}{L}\sum_{l\in[L]}B_{l}\Delta^{2}B_{l}\|_{2},$
where $\Delta:={\operatorname{diag}}(\sqrt{n_{1}},...,\sqrt{n_{K}})$.
@@
Assumption 1 requires that the communities are balanced. Assumption 2 specifies the connectivity matrices. In particular, we require the network sparsity $\rho=o(1)$. Assumption 3 requires the linear growth of the aggregated squared connectivity matrices in terms of its minimum eigenvalue, which is trivially satisfied when $B_{l,0}$’s are equal and is also used in Lei and Lin (2022). Assumption 3 combining with Assumptions 2 and 1 implies that $\sum_{l=1}^{L}B_{l}\Delta^{2}B_{l}$ is of full rank, and $Q$ is of rank $K$; see the proof of Theorem 4. Indeed, the full rank requirement could be extended to the rank-deficient case; see Zhang et al. (2022) for example. Definition 2 measures how far the local average $\sum_{l\in S_{i}}B_{l}\Delta^{2}B_{l}/s$’s deviates from their global average $\sum_{l\in[L]}B_{l}\Delta^{2}B_{l}/L$. In what follows, we call $\mathcal{H}$ the measure for the heterogeneity.
@@
Suppose Assumption 1 holds, then the output $\tilde{V}$ of Algorithm 1 satisfies
$\mathrm{dist}(\tilde{V},V)\lesssim{\max_{i\in[m]}\frac{\|\sum_{l\in S_{i}}(\overline{M}_{l}-Q_{l})/s\|_{2}^{2}}{\lambda_{{\min}}^{2}(Q)}}+{\frac{\mathcal{H}^{2}(n,\rho,L)}{\lambda_{{\min}}^{2}(Q)}}+\frac{\|\sum_{l}\overline{M}_{l}/L-Q\|_{2}}{\lambda_{{\min}}(Q)}:=\mathcal{I}_{1}+\mathcal{I}_{2}+\mathcal{I}_{3}.$
@@
The condition that $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are smaller than a small constant $c$ is a standard condition (Damle and Sun, 2020; Charisopoulos et al., 2021); it ensures that the maximum error for the local average $\max_{i\in[m]}\|\sum_{l\in S_{i}}\overline{M}_{l}/s-Q\|_{2}$ is smaller than the eigen-gap $\lambda_{{\min}}(Q)$, which further indicates that the $K$th largest eigenvalues of $\sum_{l\in S_{i}}\overline{M}_{l}/s$ are separated from its remaining eigenvalues by Weyl’s inequality. In Theorem 4, we show that this condition is satisfied by our setting under mild conditions and further provide the explicit expressions for $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$.
The results of Theorem 2 hold for any reference eigenvector $\hat{V}_{o}$ provided that we choose the true eigenspace $V$ to be aligned with $\hat{V}_{o}$ in the sense that $\operatorname{arg\,min}_{Z\in\mathcal{O}_{K}}\|\hat{V}_{o}Z-V\|_{F}=\mathbb{I}_{K}$ by noting that we have a degree to choose the version of eigenspace $V$ (Charisopoulos et al., 2021). Empirically, since it is easy to see that each local machine has the same population-wise eigenspaces, the experimental results would be similar when choosing a different sample-wise reference solution $\hat{V}_{o}$.
Remark 4
@@
The decomposed three components $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$ are explained as follows. $\mathcal{I}_{1}$ corresponds to the maximum squared error of the local estimators $\hat{V}_{i}$’s except that we use $Q$ in the denominator instead of $\sum_{l\in S_{i}}Q_{l}/s$’s; see the Davis-Kahan theorem on the perturbation of eigenvectors (Davis and Kahan (1970); see Lemma 3 in the Supplementary Materials). As a result, we only require the averaged population matrix $Q:=\sum Q_{l}/L$ be of rank $K$, and the local averaged population matrix $\sum_{l\in S_{i}}Q_{l}/s$ can be rank deficient, which is one of the benefits of combining the local eigenvectors. $\mathcal{I}_{2}$ corresponds to the effect of heterogeneity. To distinguish, we call $\mathcal{H}$ the measure for heterogeneity and call $\mathcal{I}_{2}$ the measure for effective heterogeneity. Note that $\mathcal{H}$ and $\lambda_{{\min}}(Q)$ jointly decide $\mathcal{I}_{2}$. Therefore, if the local averages are noisy from the signal such that combining them across machines could hide the true signal, i.e., $\mathcal{H}$ is large, and $\lambda_{{\min}}(Q)$ is small, then the effective heterogeneity $\mathcal{I}_{2}$ would be large. On the other hand, if the local averages are far from the signal but combining them would enhance the signal, i.e., $\mathcal{H}$ is large, but $\lambda_{{\min}}(Q)$ can also be large such that their ratio is small, then the effective heterogeneity $\mathcal{I}_{2}$ would be small. We refer to the two simulated examples in Section C of the Supplementary Materials for more illustrations on the aforementioned phenomena. $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ can be regarded as the price needed to pay for the distributed computation. Moreover, $\mathcal{I}_{3}$ corresponds to the error of the centralized estimator $\hat{V}:=\frac{1}{L}\sum_{l}\overline{M}_{l}$.
Before developing our main result, namely, the specific version of Theorem 2, we first provide Theorem 3 to study how the privacy parameters $q,q^{\prime}$ affect the upper bound for the centralized estimator $\|\sum_{l}\overline{M}_{l}/L-Q\|_{2}$. We need the following definition of $\xi(q,q^{\prime})$ and $C(q,q^{\prime})$ through out the remainder of this section to simplify the notation,
$\xi(q,q^{\prime})=\frac{1-q^{\prime}}{(q+q^{\prime}-1)^{2}}\quad{\rm and}\quad C(q,q^{\prime})=\frac{q}{(q+q^{\prime}-1)^{2}},$ (12)
where we note that when $q,q^{\prime}>1/2+\tau$ with a small constant $\tau$, $C(q,q^{\prime})$ can be regarded as a constant. We further use $\xi_{q,q^{\prime}}$ and $C_{q,q^{\prime}}$ to simplify the notation.
@@
$\small\|\overline{M}-Q\|_{2}\lesssim\max$$\bigg{\{}\frac{(C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}})^{1/2}n^{1/2}\rho\log^{1/2}(L+n)}{\sqrt{L}},$
$\quad\quad\quad\frac{(C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}})\log(L+n)}{\sqrt{L}},\,{(\rho+(q+q^{\prime}-1)\xi_{q,q^{\prime}})^{2}}\bigg{\}},$ (14)